Melanie can paint a certain barn by herself in x days. Her helper Melisa can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x = 5

Answer by Alan3354(30993)

Melanie can paint a certain barn by herself in x days. Her helper Melisa can paint the same barn by herself in 2x days. Write a rational expression for the fraction of the barn that they complete in one day by working together. Evaluate the expression for x = 5
-------------
Melanie paints 1/x of the barn per day.
Melisa paints 1/2x per day.
Together, they do 1/x + 1/2x per day
1/x + 1/2x = 3/2x per day.
-------------
At x=5:
--> 3/10 per day.
---------------
The long way for x - 5:
1/5 + 1/10 = 3/10 --> 10/3 days to paint the barn together
------------------
A shortcut for x = 5: product/sum
5*10/(5+10) = 50/15
= 10/3 days